OpenStudy (anonymous):

How do I graph x^2=4y^2? Would someone mind making an x and y chart and showing me?

OpenStudy (zzr0ck3r):

first graph \(y_1=\sqrt{\dfrac{x^2}{4}}\) and then on the same graph, graph \[y_2=-\sqrt{\dfrac{x^2}{4}}\]

OpenStudy (zzr0ck3r):

Because what you put is not a function, so we break it up into functions :)

OpenStudy (anonymous):

x^2-4y^2=0 (x-2y)(x+2y)=0 so you should draw x-2y=0 and x+2y=0 .

OpenStudy (anonymous):

look at this http://www.wolframalpha.com/input/?i=plot+x^2%3D4y^2

OpenStudy (anonymous):

Oh...how do you tell when to break up equations into two functions?

ganeshie8 (ganeshie8):

both look good to me.. amirreza's even better...

OpenStudy (zzr0ck3r):

sorry I read something completely different on my phone ....

ganeshie8 (ganeshie8):

thats zero product property @one2345 if you have an equation like : \(a*b = 0\), then it must be the case that either \(a = 0\) or \(b=0\) (or both of them equal to \(0\))

OpenStudy (anonymous):

Oh I see, thank you so much.